• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Roedelbronn

    Log entry time 04:09:26 on June18,2002

    Entry number 86325

    keyword=VDC problems

    00:28 Reset slow controls for HRS-R. When the crate came back up the VDC HV had both tripped. We tried to bring them back up to 4.0 kV but only the R_TOP was working. We tried to reset the slow controls once more and encountered the same problem.

    00:40 Took the Hall to Controled Access and entered the detector hut to try a local reset by cycling power to the power supply crate - also unsuccessful.

    01:30 Jack Segal came in and reset the power supply locally. He said our cycling of crate power should have had the same effect, and it wasn't clear why that didn't work. Jack's reset was successful, and the VDC came back up to 4 kV. He may look more closely at it during tomorrow's maintenance. He has experienced quirky behavior in resetting these types of power supplies before.